Week 4

1. This artist's conception imagines a 'wheel-shaped' space station. The idea is that the 'wheel' would turn on its axis at a rate such that the acceleration in the rim of the wheel would be 9.8 m/s^2. (This would make inhabitants feel as though they had their normal weight.) Suppose that the radius of the wheel is 1.0 x 10^2 m. What is the required period of rotation of the space station about its axis? (Follow the significant number for the final answer).
1
Expert's answer
2020-03-02T10:16:28-0500

As per the question,

Acceleration in the rim of the wheel is "(a)=9.8 m\/sec^2"

Radius of the wheel (R)="1.0 \times 10^2 m"

We know that "a=R\omega^2"

"\omega=" angular velocity of the wheel

We know that angular velocity "(\omega)=\dfrac{2\pi}{T}"

where T is the time period.

So,

"9.8 m/sec^2=1.0\times \dfrac{2\pi}{T}"

"\Rightarrow T=\dfrac{2\pi}{9.8}=0.64 sec"

or 0.6 sec